# Uncertainty quantification in fine-tuned LLMs using LoRA ensembles

### v1

@@ 5.3 Prior Selection
An inappropriate selection of the prior can lead to a posterior that performs poorly on the task at hand. Consequently, we adjust the variance of the prior with the aim of optimizing the posterior quality, as measured by the log likelihood on the validation data. Different choice for the priors variance results in adjusting L2 regularization loss according to
@@ 6 Uncertainty quantification
We quantify total predictive uncertainty, with both aleatoric and epistemic contributions, using predictive entropy
@@
where $s^{*}$ is the input string, $\mathcal{D}$ is the training data, and $c$ represents all possible output tokens. The output probabilities $p(c|s^{*},\mathcal{D})$ are computed by averaging softmax outputs from network parameters $\theta$, sampled from the Bayesian posterior $p(\theta|\mathcal{D})$.
We quantify epistemic uncertainty by the average mutual Shannon information between the model parameters and a test data sample, given the training dataset (Mackay, 1992)
$\begin{split}&\text{MI}(\theta,t^{*}|s^{*},\mathcal{D})\\ &=H(\theta|\mathcal{D})-\mathbb{E}_{t^{*}\sim p(t^{*}|s^{*},\mathcal{D})}\left[H(\theta|t^{*},s^{*},\mathcal{D})\right]\\ &=H(t^{*}|s^{*},\mathcal{D})-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}\left[H(t^{*}|s^{*},\theta)\right].\\ \end{split}$ (5)
The mutual information MI in Eq. (5) is the expected change in the posterior parameter distribution when a new data point is added to the training set (first row of RHS) (Houlsby et al., 2011), and can be conveniently calculated in terms of the entropy of the predictive distributions (second row of RHS).
Even though entropic uncertainty measures have been shown to contain valuable information for neural network uncertainty quantification (Linander et al., 2023), one should be aware of unintuitive properties shown in recent studies (Wimmer et al., 2023).
@@ 7.1 Datasets
We evaluate our methods on three multiple-choice QA datasets: CommonsenseQA (CQA) (Talmor et al., 2019), and the Social Sciences (SS) as well as STEM components from MMLU (Hendrycks et al., 2021). Detailed information about the training and validation (test) splits is available in Appendix A.
@@ 7.2 LoRA Ensembles
We train LoRA ensembles using the pre-trained 7 billion parameter Mistral-7b model (Jiang et al., 2023). The LoRA parametrization was implemented using PEFT (Mangrulkar et al., 2022). Our ensembles contain 5 members $(M=5)$, but we also investigate larger ensembles in our ablation studies. Adhering to the method of Hu et al. (Hu et al., 2021), we applied the adapter $\Delta W=\alpha BA$ exclusively to the query and value matrices within the self-attention modules of Mistral-7b, with $\alpha$ set to 32. The rank was set to $r=8$, which gives $3,407,872$ trainable parameters. The adapter matrices $B$ are initialized to zero, while the entries of $A$ are initialized using a Kaiming-uniform distribution (He et al., 2015).
The ensemble members are trained using the prescribed regularisation from Section 5.3, corresponding to using the pretrained Mistral-7b model as a prior. We use the Adam optimizer, training for 10 epochs with learning step sizes of $5\cdot 10^{-6}$ and $1.25\cdot 10^{-6}$, and batch sizes of 8 and 2 for the CQA and MMLU datasets, respectively. These learning rates are smaller compared to those in (Wang et al., 2023) to better refine the learning curve before the models become over-fitted. The token size limit is fixed at 128 for CQA and 512 for MMLU datasets. We performed an ablation study on L2 regularization $\lambda/2=0.01,0.1,1,10$, finding optimal performance at $\lambda/2=1$ and $10$ for the CQA and MMLU datasets, respectively. The training focuses solely on the single token output representing the answer, with all other token outputs being masked.
@@
Figure 1: Performance of the LoRA ensembles trained and evaluated on either CQA, MMLU STEM, or MMLU SS dataset. The metrics include accuracy, log likelihood loss, and expected calibration error (ECE). The number of ensemble members is either $M=5$ or $M=1$. For a LoRA ensemble with $M=1$ the average results over 5 distinct realizations are shown.
@@ 7.3 Performance metrics
In Fig. 1, we illustrate the performance metrics for LoRA ensembles with both $M=1$ and $M=5$ members. The ensembles are trained on multiple-choice QAs from CQA (first column), MMLU STEM (second column), and MMLU Social Studies datasets (third column).
@@
The ensemble members start over-fitting after a couple of epochs as can be seen by the validation losses in the second row of panels, but interestingly without considerable drop in accuracy in the first row of panels. As will be seen in more detail in Section 7.5, this signals that the ensemble members become more overconfident on predictions that are wrong, corroborated by the increasing expected calibration error with epoch seen in the bottom row of Fig. 1. This is consistent with the recent reports that fine-tuned LLMs often exhibit overconfidence (Jiang et al., 2021; Lin et al., 2022; Xiao et al., 2022b; He et al., 2023; Tian et al., 2023; OpenAI et al., 2023).
@@
For fine-tuning on CQA seen in the first column, the overfitting is significantly reduced for ensembles compared to individual members, as can be seen by the gap between the loss curves for $M=1$ and $M=5$ for later epochs. This means that ensemble members output more confident, but different predictions. This is a common feature of Bayesian posteriors (Blundell et al., 2015; Zhang et al., 2020; Kristiadi et al., 2020; Ober & Aitchison, 2021; Fortuin et al., 2022; Aitchison et al., 2021; Yang et al., 2024). This signals high epistemic uncertainty in this regime, indicating that the validation set is perceived as out-of-domain. There is no significant difference between single member and ensemble metrics for models trained on the MMLU datasets. We attribute this to the fact that the MMLU training datasets are small, see Appendix A, facilitating fast overfitting before any significant generalization is attained.
Figure 2: AUROC computed for the CQA, MMLU STEM, and MMLU Social Studies multiple-choice QA datasets. In each subpanel, one dataset is used for fine-tuning (T) on the training subset and two datasets for the AUROC evaluation (V) on the validation subset. The LoRA ensemble size is $M=5$.
7.4 Uncertainty metrics visualized by AUROC
Area under receiver operator curve (AUROC) is commonly used as a metric for out-of-domain detection (Hendrycks & Gimpel, 2016), but can also be used more generally to measure how well a model can separate two datasets. By evaluating AUROC using uncertainty measures, we can gain understanding of which of the datasets is more consistently perceived by the model as more confusing (using entropy) or as more out-of-domain (using mutual information).
In Fig. 2, we show the AUROC results comparing in- and out-of-domain validation datasets. The ensembles were fine-tuned on the training subset of the CQA, MMLU STEM, and MMLU Social Studies datasets. For the AUROC evaluation, we use the validation subset of these datasets as in-domain samples, and the validation subset of other datasets as out-of-domain.
By analyzing the AUROC curves for the ensemble trained on CQA in the first column of Fig. 2, it is evident that the model consistently perceives samples from CQA as more in-domain compared to those from MMLU datasets, as indicated by an AUROC greater than 0.5. This aligns with expectations. However, the AUROC is not close to 1, suggesting that a significant portion of the MMLU dataset is perceived similarly to CQA in terms of entropy and mutual information, with MMLU Social Sciences appearing closer than MMLU STEM as can be seen by the lower AUROC values. A somewhat similar pattern is observed for the ensemble trained on MMLU Social Sciences.
For the ensemble trained on MMLU STEM in the second column, a peculiar behavior is observed, where the AUROC is below 0.5. We interpret this as follows: MMLU STEM poses greater complexity for the LLM architecture, hence, even with training, it struggles to identify the right features for correct answers. Conversely, the ensemble retains robust knowledge about features pertinent to MMLU SS and CQA. A mutual information score below 0.5 indicates that the ensemble perceives the validation set of MMLU STEM as more out-of-domain compared to the other datasets, even though the model was fine-tuned on the training subset of MMLU STEM.
@@ 7.5 Dynamics of uncertainty metrics
To analyze the trends observed through the AUROC curves in more detail, we also produce data-resolved dynamics of uncertainty measures by looking at two-dimensional histogram plots of mutual information and entropy(Linander et al., 2023).
@@
In Fig. 3, we show the epoch evolution of uncertainty measures for a LoRA ensemble that was fine-tuned and evaluated on the CQA dataset. The mutual information (MI) on the vertical axis, and the predictive entropy (Entropy) on the horisontal axis, are binned over every sample of the validation set of CQA and the number of samples in each bin is indicated by the colorbar. We distinguish between incorrect and correct predictions to enhance interpretability. The ensemble attains minimal validation loss at epoch 3, as shown in Fig. 1. Epochs 1 and 6 exhibit underfitting and overfitting, respectively. The true posterior $p(\theta|D)$ should yield nearly zero mutual information when evaluated on the validation dataset, which should be in-domain by design. Samples with high mutual information indicate that parts of the target domain are perceived as out-of-domain by the model, showing the most important regions to consider for improving the training dataset targetting a better posterior approximation. We observe a moderate spread in entropy on the horizontal axis, indicating that the model predicts a broad distribution even though the sample is considered in-domain through their small mutual information. This spread can be attributed to the questions where the model does not have enough context or expressive power for a distinct answer.
@@
Fig. 4 shows the uncertainty histograms for MMLU STEM and MMLU SS, calculated using an ensemble trained on CQA. Starting with MMLU STEM in the first and second row of panels, we draw two conclusions. First, out of the wrongly classified samples for the optimal epoch 3 shown in the second row, second column, there is a large fraction with high predictive entropy and low mutual information compared to the more moderate spread of predictive entropy in the corresponding panel for MMLU SS in the last row, second column. This affirms conclusions from Section 7.4 regarding the relatively higher complexity of STEM compared to SS. The ensemble trained on CQA perceives the STEM samples equally in-domain as the SS samples, as evident by the similar mutual information, but predict a broader distribution indicating a lack of knowledge. Secondly, in the over-fitting regime corresponding to the last column, the mutual information of the wrongly classified samples of the second row increase more than the correctly classified samples in the first row. This indicates that the ensemble forgets the samples with high predictive entropy quicker, i.e. it forgets the confusing samples faster.
@@
Figure 3: Histograms of predictive entropy and mutual information for a LoRA ensemble trained and evaluated on the CQA dataset. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. This figure illustrates the evolution of uncertainty measures for in-domain data across training epochs, differentiated by correct and incorrect predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
@@
Figure 4: Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA and evaluated on MMLU STEM and MMLU SS datasets. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. This panels illustrates the evolution (left to right) of uncertainty measures for out-of-training-domain data across training epochs, differentiated by correct and incorrect predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values.
Continuing with MMLU SS in the third and fourth row, we observe a decrease in predictive entropy in the over-fitting regime, indicating that the ensemble does find features giving confident but wrong predictions.
@@
Interestingly, the evolution of entropic quantities shows a quantitative difference: For MMLU STEM, there is a significant density of high-entropy, low-mutual-information samples at epoch 6, whereas for MMLU SS, the data with high entropy is more dispersed in terms of mutual information. This suggests that the fine-tuned model perceives confusing samples from MMLU STEM and MMLU SS differently, forgetting the confusing samples from MMLU SS more rapidly.

### v2

@@ 5.3 Prior Selection
An inappropriate selection of the prior can lead to a posterior that performs poorly on the task at hand. Consequently, we adjust the variance of the prior with the aim of optimizing the posterior quality, as measured by the log likelihood on the validation data. Different choice for the prior variance results in adjusting L2 regularization loss Hoffmann & Elster (2021); Balabanov et al. (2023). In the case of LoRA ensembles, this loss is given by
@@
5.4 Posterior Quality of LoRA Ensembles
LoRA deep ensembles approximate the true posterior over model parameters for the fine-tuning task via variational inference, and the trustworthiness of the resulting uncertainty estimates depends on how closely this approximation matches $p(\theta|\mathcal{D}_{\text{fine-tune}})$. Hamiltonian Monte Carlo (HMC) remains the gold standard for posterior estimation because it generates nearly exact samples and thus yields the most faithful uncertainty estimates. In principle, one would evaluate any approximation by measuring its agreement and variance relative to the HMC posterior. In practice, HMC-derived posteriors do not always correlate with the best predictive performance: they can be poorly calibrated and sensitive to distributional shifts Izmailov et al. (2021), and stochastic ensembles can outperform them in accuracy and NLL Balabanov et al. (2023). This does not diminish their status as the best available posterior approximation, and as such they are the target for any potential approximation method. This closeness can be quantified using agreement and variance metrics Izmailov et al. (2021). Unfortunately, running HMC on large language models and datasets is prohibitively expensive, placing it beyond our computational budget and preventing direct comparison.
Although direct comparison is not possible, prior work has shown that multimodal methods such as deep ensembles yield accurate posterior approximations and significantly outperform single-mode approaches Fort et al. (2019a); Izmailov et al. (2021); Balabanov et al. (2023). Building on these insights, we argue that our LoRA ensemble–based method delivers high-quality posterior estimates under realistic training and inference constraints, while also surpassing single-mode baselines such as the Laplace approximation, MC dropout, and related techniques. We acknowledge that Yang et al. Yang et al. (2024) examined LoRA-tuned LLMs primarily through a single-mode Laplace approximation, comparing it against several other methods—including a small ensemble—and demonstrated its performance benefits. Although their results are pertinent to performance-centric applications, they do not address the uncertainty quantification focus of our study. In fact, the Laplace method’s apparent robustness to distribution shifts and “good” calibration may actually indicate a poor approximation of the true posterior Izmailov et al. (2021).
@@ 6 Uncertainty quantification
6.1 Epistemic model and aleatoric data uncertainty
Uncertainty quantification is an important topic that assesses the reliability of model predictions (Gal & Ghahramani, 2016). In general, the total uncertainty in a model’s prediction can be attributed to multiple sources. One widely used categorization is the decomposition into epistemic (or model) uncertainty—arising from insufficient knowledge or data to fully learn the underlying distribution—and aleatoric (or data) uncertainty—stemming from the inherent noise and randomness in the data generation process.
An example of aleatoric data uncertainty is given by the input sequence “There is a single item in the box. The color of that item is ”. One would expect a model trained on a large corpus of text to produce a predictive distribution with support on the different colors. This posterior distribution tells us that the model is fairly certain that the next token is a color, but that there is an uncertainty about which color. The uncertainty about the color cannot be decreased by improving the training data; it is inherent in the input sequence.
On the other hand, the input sequence “Large language models are ” presents a simple conditional completion task, but we would not trust the output predictions from a model trained on a text corpus compiled before the concept was introduced. This epistemic model uncertainty is not captured in the output distribution, but rather in the uncertainty of the model parameters themselves. This model uncertainty can be systematically calculated within a Bayesian formalism, where it corresponds to the shape of the posterior distribution of model parameters conditioned on the training data.
6.2 Entropic uncertainty metrics
Depending on the approach used to estimate uncertainty, existing methods can be broadly divided into intrinsic and extrinsic methods. Extrinsic methods involve post-hoc analysis with auxiliary models to estimate uncertainty Lahlou et al. (2023); Kristiadi et al. (2021). In contrast, intrinsic methods incorporate uncertainty estimation directly within the model. Here, we employ a common intrinsic methodology based on entropic metrics Malinin & Gales (2021) to evaluate uncertainty within a Bayesian formulation, applying it to our LoRA ensemble approximation of the Bayesian posterior for fine-tuned models.
The total predictive uncertainty, encompassing both aleatoric data and epistemic model components, can be quantified using the total predictive entropy:
@@ 6.2 Entropic uncertainty metrics
where $s^{*}$ is the input string, $\mathcal{D}$ is the training dataset, and $c$ denotes all possible output tokens. The output probabilities $p(c\mid s^{*},\mathcal{D})$ are computed by averaging softmax outputs across model parameters $\theta$ sampled from the Bayesian posterior $p(\theta\mid\mathcal{D})$.
Epistemic model uncertainty is defined as the variability in predictions that arises from sampling model parameters $\theta$ from the posterior distribution $p(\theta\mid\mathcal{D})$. This variability captures how the distribution of predicted outputs shifts with each parameter set, weighted by its posterior probability. A common metric for this purpose is the average mutual information between the model parameters and output, conditioned on the training data Mackay (1992); Malinin & Gales (2021):
$\begin{split}&\text{MI}(\theta,t^{*}|s^{*},\mathcal{D})\\ &=H(t^{*}|s^{*},\mathcal{D})-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}\left[H(t^{*}|s^{*},\theta)\right].\\ \end{split}$ (5)
For ensemble posteriors, zero mutual information implies that all ensemble members yield the same predictive distribution, indicating no variability in the outputs and thus zero epistemic uncertainty. On the contrary, if the mutual information is high, there is significant variability in the predictive distribution across the ensemble, reflecting high epistemic uncertainty.
Using these metrics, we can identify three regimes that describe how the model perceives data:
  • Known-certain: Low total and epistemic uncertainty; the posterior ensemble strongly agrees on a narrow, confident prediction distribution.
  • Known-uncertain: High total uncertainty, low epistemic uncertainty; uncertainty arises from data noise or model expressiveness limitations, with the ensemble members agreeing on a broad distribution.
  • Unknown: High total and epistemic uncertainty; the posterior members disagree on the prediction distribution, highlighting a lack of knowledge.
In this work, we use uncertainty quantification from entropic metrics of the LoRA ensemble posterior approximation to better understand how fine-tuning reshapes the knowledge of the model, revealing shifts in data characteristics and how the model adapts to new information. In particular, we study the dynamics of the uncertainty metrics during fine-tuning.
@@ 7.1 Datasets
We evaluate our methods on three multiple-choice QA datasets: CommonsenseQA (CQA) Talmor et al. (2019), and the Social Sciences (SS) as well as STEM components from MMLU Hendrycks et al. (2021). The ensemble is trained on the CQA training dataset but tested on three different validation datasets: the (in-domain) CQA validation dataset and the (out-of-domain) MMLU STEM and MMLU SS validation datasets. Detailed information about the training and validation (test) splits is available in Appendix A.
@@ 7.2 LoRA Ensembles
We train LoRA ensembles using the pre-trained 7-billion-parameter Mistral-7B model Jiang et al. (2023). The LoRA parametrization is implemented using PEFT Mangrulkar et al. (2022). We evaluate ensembles with sizes $M\in\{1,5,10,20\}$ in our experiments and ablation studies. Following the method of Hu et al. Hu et al. (2021), we apply the adapter $\Delta W=\alpha BA$ exclusively to the query and value matrices within the self-attention modules of Mistral-7B, with $\alpha=32$. The adapter rank is set to $r=8$, resulting in $3{,}407{,}872$ trainable parameters. The adapter matrices $B$ are initialized to zero, and the entries of $A$ are initialized using a Kaiming-uniform distribution He et al. (2015).
The ensemble members are trained using the prescribed regularisation from Section 5.3, corresponding to using the pretrained Mistral-7b model as a prior. We use the Adam optimizer, training for 10 epochs with learning step sizes of $5\cdot 10^{-6}$ and batch sizes of 8. This learning rate is smaller compared to those in Wang et al. (2023) to better refine the learning curve before the models become overfitted. The token size limit is fixed at 128 for CQA and 512 for MMLU. We performed an ablation study on L2 regularization $\lambda/2=0.01,0.1,1,10$, finding optimal performance at $\lambda/2=1$. The training focuses solely on the single token output representing the answer, with all other token outputs being masked.
@@ 7.3 Performance metrics
In Fig. 1, we present the performance metrics for LoRA ensembles with ensemble sizes $M=1$, $5$, $10$, and $20$. The ensembles are trained on multiple-choice questions from the CQA training dataset and evaluated on three validation datasets: CQA (first column), MMLU STEM (second column), and MMLU Social Studies (third column).
As shown in Fig. 1, the loss curves for ensemble sizes $M=5$, $10$, and $20$ converge closely, indicating that small ensembles ($M=5$) provide posterior approximations comparable to those of larger ensembles ($M=20$). This convergence is particularly interesting and practically relevant, as it suggests that LoRA ensembles can achieve accurate uncertainty estimates without significantly increasing inference costs. Based on this observation, we use $M=5$ ensembles for uncertainty quantification in the following section.
Figure 1: Performance of LoRA ensembles trained on CQA and evaluated on the CQA, MMLU STEM, and MMLU SS datasets. Metrics (computed over the ensemble mean distribution) include accuracy, negative log-likelihood (NLL) loss, and expected calibration error (ECE). We show ensemble sizes $M\in\{1,5,10,20\}$; for $M=1,5,10$, results are averaged over 20, 4, and 2 independent realizations, respectively, while $M=20$ uses a single run. Metric values at epoch 1 (underfitting), epoch 3 (optimal), and epoch 6 (overfitting) are highlighted.
@@
The ensemble members start over-fitting after a couple of epochs as can be seen by the validation loss in the second row of panels, but interestingly without considerable drop in accuracy in the first row of panels. As will be seen in more detail in Section 7.4, this signals that the ensemble members become more overconfident on predictions that are wrong, corroborated by the increasing expected calibration error with epoch seen in the bottom row of Fig. 1. This is consistent with the recent reports that fine-tuned LLMs often exhibit overconfidence Jiang et al. (2021); Lin et al. (2022); Xiao et al. (2022b); He et al. (2023); Tian et al. (2023); OpenAI et al. (2023).
@@
For fine-tuning on CQA, shown in the first column, overfitting is significantly reduced for ensembles compared to individual models, as evidenced by the gap between the loss curves for $M=1$ and those for $M=5,10,20$ at later epochs. This suggests that ensemble members produce more confident yet diverse predictions. This means that ensemble members output more confident, but different predictions. This is a common feature of Bayesian posteriors Blundell et al. (2015); Zhang et al. (2020); Kristiadi et al. (2020); Ober & Aitchison (2021); Fortuin et al. (2022); Aitchison et al. (2021); Yang et al. (2024). This signals high epistemic uncertainty in this regime, indicating that the validation set is perceived as “unknown” to the model.
@@ 7.4 Dynamics of uncertainty metrics
We analyze data-resolved dynamics of uncertainty measures using two-dimensional histogram plots of mutual information and entropy Linander et al. (2023). This approach allows us to effectively disentangle different sources of uncertainty, providing a detailed view of how the model perceives data across the “known-certain,” “known-uncertain,” and “unknown” regimes during fine-tuning. By visualizing the relationship between mutual information and entropy, we can track how the model’s knowledge evolves, identify patterns in data complexity, and assess the model’s ability to adapt to fed information. This method not only highlights the dynamics of learning, but also helps identify limitations of the model, offering actionable insights for further improvement.
@@
In Fig. 2, we show the epoch evolution of uncertainty measures for a LoRA ensemble that was fine-tuned and evaluated on the CQA dataset. The mutual information (MI) on the vertical axis, and the predictive entropy (Entropy) on the horisontal axis, are binned over every sample of the validation set of CQA and the number of samples in each bin is indicated by the colorbar. We distinguish between incorrect and correct predictions to enhance interpretability. The ensemble attains optimal validation loss at epoch 3, as shown in Fig. 1. Epochs 1 and 6 exhibit underfitting and overfitting, respectively. The true posterior $p(\theta|D)$ should yield nearly zero mutual information when evaluated on the validation dataset, which should belong to “known” data by design. However, in each of the underfitted, optimal, and overfitted regimes, there is a fraction of validation samples that exhibit a moderate spread in mutual information. Samples with high mutual information indicate that parts of the target domain are perceived as “unknown” by the ensemble, highlighting regions where expanding or refining the training data could most improve performance.
By examining the mean and median of the entropic metrics in Fig. 2, we can observe overall patterns in the validation data perception by the ensemble. Interestingly, the ensemble becomes progressively more confident on correct QAs, reaching its highest confidence once the ensemble overfits, where most correct QAs are predicted with almost zero uncertainty. Although overall performance (as measured by loss shown in Fig. 1) degrades at epoch 6 compared to epoch 3, the model has actually learned the bulk of the validation dataset more effectively. The loss increase arises because the ensemble has also become more confident about incorrect QAs, increasing the cross-entropy. However, almost all of these incorrect predictions are marked by high mutual information, placing them in the “unknown” regime. Since these points can be identified by their high mutual information, they can be filtered out at inference time or used to refine the training procedure (e.g., by including more instances of these data). This result is noteworthy because it shows that the ensemble continues to improve even after reaching its minimum loss at epoch 3. By epoch 6, it performs better on most of the dataset, although its increased confidence in the remaining poorly performing data inflates the overall loss. Nonetheless, these problematic data points remain mostly distinguished by high epistemic uncertainty.
@@
Fig. 3 shows the uncertainty histograms for MMLU STEM and MMLU SS, calculated using an ensemble trained on CQA. Starting with MMLU STEM in the first and second row of panels, we draw two key insights. First, among the wrongly classified samples for the optimal epoch 3 (second row, second column), a large fraction falls into the “known-uncertain“ regime, characterized by high predictive entropy but low mutual information. This is in contrast to MMLU SS, which exhibits a more moderate spread of predictive entropy in the corresponding panel (last row, second column). This observation highlights the relatively higher complexity of STEM compared to SS datasets. The ensemble perceives the STEM samples as equally “known” as SS samples (similar mutual information) but assigns them broader predictive distributions, indicating greater “uncertain” characteristics due to a lack of precise knowledge.
@@
Figure 2: In-distribution evolution of uncertainty during fine-tuning. Histograms of predictive entropy and mutual information for a LoRA ensemble trained and evaluated on the CQA dataset. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. This figure illustrates the evolution of uncertainty measures for in-domain data across training epochs (columns), differentiated by correct (top row) and incorrect (bottom row) predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values. Color represents the number count in each bin.
@@
Figure 3: Out-of-distribution evolution of uncertainty during fine-tuning. Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA and evaluated on MMLU STEM (top rows) and MMLU SS (bottom rows) datasets. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. These panels illustrates the evolution of uncertainty measures for out-of-training-domain data across training epochs (columns left to right), differentiated by correct (first and third row) and incorrect (second and forth row) predictions. We also depict the corresponding mean (red) and median (green) entropy and mutual information values. Color represents the number count in each bin.
Figure 4: Late overfitting regime (epoch 10). Histograms of predictive entropy (Entropy) and mutual information (MI) for a LoRA ensemble trained on CQA (left column) and evaluated on MMLU SS (middle column) and MMLU STEM (right column) datasets. The ensemble consists of $M=5$ members, with a LoRA rank of 8, $\alpha=32$, and an L2 LoRA loss of 1. There is still a significant fraction of samples (lower right panel) in MMLU STEM with low mutual information and high total uncertainty.
Second, continuing with MMLU SS in the third and fourth rows of Fig. 3, we observe a significant decrease in predictive entropy as the number of epochs increases, indicating that the ensemble learns features that enable it to classify samples with greater confidence. This behavior is similar to CQA in Fig. 2, but the key difference is that, in this case, the data is from an out-of-domain dataset. This demonstrates that the ensemble continues to improve on the majority of MMLU SS samples, despite not being specifically trained on them, suggesting a similarity between the CQA and MMLU SS domains.
@@
Interestingly, the evolution of entropic quantities reveals a key difference between the datasets. For MMLU STEM, there is a notable density of “known-uncertain” samples (high entropy, low mutual information) (see Fig. 3, second row), whereas for MMLU SS, high-entropy samples are more evenly distributed in terms of mutual information. The density of “known-uncertain” MMLU STEM samples largely persists despite overfitting (see Fig. 4), suggesting that the fine-tuned model ensemble perceives confusing samples from MMLU STEM as being more consistently “known-uncertain,” whereas confusing samples from MMLU SS are forgotten more rapidly, transitioning into the “unknown” regime.